The students on the Plaid team sold cookies to raise money. Peanut butter cookies sold best. The students sold half as many chocolate chip cookies as peanut butter cookies. Twenty dozen cookies were sold all together. How many cookies of each type were sold?
Answer by Aswathy(23)

Let the number of chocochip cookies sold be x.
As the number of peanut butter cookies sold is half the number of chocochip cookies sold by them and it is also given that the sum of both the cookies is 20 dozen.
So according to the question,
x+(x/2)=20*12
(3x/2)=240
3x=480
x=160
(x/2)=80